# Transmission lines; Defect detection; UAV imagery; Power industry; Transformer-based detectors

### v4

@@ Abstract
Transmission lines; Defect detection; UAV imagery; Power industry; Transformer-based detectors
@@ 1 INTRODUCTION
With the rapid advancement of smart grid development, transmission line defect detection has increasingly adopted deep learning-based object detection methodsLi et al. (2023). Alongside this progress, numerous general-purpose detectors—such as YOLO(Redmon et al., 2016), DINOVaswani et al. (2017), and RT-DETR(Zhao et al., 2024)—have been developed. While these models achieve remarkable success in generic scenarios, most of them on the COCO datasetLin et al. (2014), they face notable limitations in transmission line defect detection under complex conditions involving occlusion, background clutter, and variable lightingLiu et al. (2023). Common issues include missed detections, false positives, and false negatives, which substantially undermine the practical utility of UAV-acquired imagery in real-world transmission line maintenanceDeng et al. (2024). These shortcomings are consistently reported in previous studies and further confirmed by our preliminary experiments, highlighting the critical need for specialized detection frameworks capable of accurately identifying small-scale defects in UAV operational scenarios.
@@
    We propose a lightweight attention block with a dual-domain Frequency–Channel–Spatial Aggregation (FCSA) mechanism, integrated in a cross-stage partial manner. This design jointly enhances global semantic reasoning and local detail preservation, enabling the network to better capture anisotropic defect patterns such as missing bolts, loose clamps, and sagging wires.
@@ 3.1 Overview
In this study, we employ RE-DETR-r18Zhao et al. (2024) as the baseline detection framework owing to its superior accuracy and robustness in complex scenarios. Traditional DETR models often suffer from slow convergence and substantial computational overhead, which hinder their deployment in UAV-based inspection tasks.
Addressing these inherent limitations, RE-DETR incorporates an enhanced encoder–decoder architecture together with refined query interaction mechanisms, while for the backbone network, RT-DETR-r18 adopts a CNN-based design, namely the popular ResNet-18He et al. (2016). From the perspective of real-time performance, employing a CNN architecture for feature extraction further enhances the real-time capability of DETR-like models, thereby improving their practical applicability. As a result, RE-DETR not only accelerates convergence and improves detection precision, but also preserves the global context modeling capability of Transformer-based frameworks(Vaswani et al., 2017).In addition, its lightweight optimization strategies—specifically, the design choice to apply the encoder only to high-level semantic features (S5) with relatively small spatial dimensions—effectively reduce resource consumption. This balances performance and speed by significantly decreasing computational overhead and accelerating inference without compromising accuracy, thereby making RE-DETR particularly suitable for post-flight UAV imagery analysis, where large volumes of data must be processed efficiently under computational and energy constraints(Tian et al., 2025).
@@
First, the backbone adopts the Edge-Enhanced ResNet (EE-ResNet) structure to strengthen low-level detail preservation. Second, a compact Space-to-Depth Convolution (S2DConv) module is inserted before each downsampling layer to reorganize spatial features in a manner that preserves critical spatial information while reducing resolution, thereby improving the recall rate of extremely small objects. Third, a lightweight Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM) is embedded into the backbone network to enhance directional sensitivity and fine-grained structural modeling, effectively capturing anisotropic features of typical defects such as missing bolts, loose clamps, or sagging wires. Finally, a Focaler-Wise-SIoU loss function is employed to jointly optimize classification and localization. This loss function adaptively balances easy and hard samples while providing more accurate bounding box regression through shape-aware optimization, which further improves detection robustness for small and ambiguous objects. Detailed descriptions of the proposed modules can be found in Sections 3.2 to  3.5.
@@ 3.2 Edge Enhanced ResNet
Thus, during deployment, EECConv degenerates into a single $3\times 3$ convolution followed by activation:
@@ 4.2 Model Training and Evaluation Metrics
The proposed model is implemented based on the PyTorch deep learning framework and is built upon the RE-DETR architecture released by Ultralytics. Training is conducted on a high-performance workstation equipped with NVIDIA GeForce RTX 3090 GPU with 24 GB of memory. The software environment consists of Ubuntu 20.04, Python 3.10.14, PyTorch 2.3.1, and CUDA 12.1, enabling efficient GPU acceleration. The model is trained for 100 epochs using the AdamW optimizer.
@@ 4.4 Ablation Study of the Proposed Method
$\checkmark$ $\times$ $\checkmark$ $\times$ $72.8$ $0.458$ $0.166$ $0.165$ $0.072$
$\checkmark$ $\times$ $\times$ $\checkmark$ $57.0$ $0.500$ $0.204$ $0.207$ $0.082$
@@
Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM) Adding the Cross-Stage Dual-Domain Multi-Scale Attention Module results in a slight decrease in performance. Specifically, precision drops from 0.486 to 0.458, and recall decreases from 0.195 to 0.166. The mAP50 metric also drops from 0.190 to 0.165, indicating that although this module may help suppress complex background interference, it slightly compromises the model’s accuracy in this configuration. Given that transmission line defects are typically small and can be visually close to background features, this module’s background suppression may sometimes inadvertently impact the model’s ability to detect small-scale defects.
@@ 4.5 Comparisons With Previous Methods
Table LABEL:tab:model_performance summarizes the overall performance of representative detectors on the CSG-ADCD dataset. The comparison covers lightweight and standard variants of YOLO (v5, v8, v10, v11), Transformer-based models such as RT-DETR and Conditional-DETR, as well as classical baselines including RetinaNet and DETR. Key indicators, including the number of parameters, computational complexity (GFLOPs), Precision, Recall, $\text{mAP}_{50}$, and $\text{mAP}_{50:95}$, are reported to provide a comprehensive view of accuracy–efficiency trade-offs. The best values in each column are highlighted in bold for clarity.
@@
Table 3: Model Performance Evaluation: Precision, Recall, and mAP for CSG-ADCD (best results highlighted in bold).
@@
YOLO v5n 2.2 5.8 0.707 0.131 0.131 0.0628
YOLO v5s 9.1 23.8 0.346 0.148 0.140 0.0646
YOLO v5m 25.1 64.0 0.548 0.164 0.162 0.0717
YOLO v8n 2.7 6.8 0.580 0.134 0.133 0.0613
YOLO v8s 11.1 28.5 0.375 0.155 0.147 0.0661
YOLO v8m 23.2 67.5 0.403 0.169 0.163 0.0719
YOLO 10n(Wang et al., 2025) 2.7 8.2 0.707 0.125 0.133 0.0627
YOLO 10s 8.0 24.5 0.536 0.144 0.157 0.0725
YOLO 10m 16.6 63.5 0.386 0.170 0.162 0.0753
YOLO 11n(Khanam and Hussain, 2024) 2.6 6.3 0.638 0.135 0.136 0.0641
YOLO 11s 9.4 21.3 0.516 0.166 0.155 0.0717
YOLO 11m 20.0 67.7 0.399 0.181 0.173 0.0776
RT-DETR-l 32.8 108.0 0.480 0.189 0.199 0.0890
RT-DETR-x 65.5 222.5 0.485 0.199 0.216 0.0990
RT-DETR-R18 19.8 57.0 0.369 0.177 0.163 0.0738
RT-DETR-R34 31.2 88.8 0.354 0.127 0.117 0.0533
RT-DETR-R50 42.9 134.8 0.363 0.120 0.115 0.0514
RT-DETR-R101 74.7 247.1 0.467 0.185 0.193 0.0879
Conditional-DETR-R18(Meng et al., 2023) 0.309 0.068 0.038 0.0110
Conditional-DETR-R34 0.447 0.082 0.034 0.0090
Conditional-DETR-R50 44.0 89.5 0.568 0.159 0.172 0.0660
DETR-R50(Carion et al., 2020) 0.301 0.138 0.124 0.0430
RetinaNet(Lin et al., 2018) 12.0 33.8 0.394 0.029 0.013 0.0040
DINO(Zhang et al., 2022) 0.496 0.257 0.224 0.0986
FFCA-YOLO(Zhang et al., 2024) 7.1 51.4 0.680 0.130 0.142 0.0596
@@
To further highlight performance on small objects and defect-specific categories, Table LABEL:tab:model_performance2 reports additional evaluation metrics. These include Recall and $\text{mAP}_{50}$ for Polluted Insulator (jyzwh) and Damaged Pole (dgss), alongside scale-sensitive measures $\text{AP}_{s}$ and $\text{AP}_{m}$. Such results enable a fine-grained comparison of detection capabilities under challenging UAV inspection scenarios.
@@ 4.6 Generalization Experiments for the Proposed Methods
Since small objects account for a large proportion in this dataset, to further validate the generalization capability of the proposed method, we additionally conduct generalization experiments on the VisDrone(Zhu et al., 2022) dataset. VisDrone is a large-scale benchmark dataset specifically constructed for visual perception tasks under the perspective of unmanned aerial vehicles (UAVs). Created by the Aerial Vision Group at the University of Science and Technology of China, the dataset comprises over 26,000 high-resolution images and 263 video sequences, captured across diverse urban and suburban environments such as streets, rooftops, playgrounds, and traffic scenes. It includes more than 100,000 manually annotated object instances spanning 10 common categories, including pedestrians, cars, buses, trucks, bicycles, and others. Each object is labeled with rich metadata, including bounding boxes, object class, occlusion level, and truncation ratio. Due to its dense object distribution, high proportion of small objects, severe occlusion, and large-scale variation in viewpoint and illumination, VisDrone presents considerable challenges for detection and tracking models. Consequently, it has been widely adopted as a standard benchmark to evaluate the robustness and generalization ability of UAV-based visual recognition algorithms in real-world complex scenarios. To further assess the generalization capability of the proposed TinyDef-DETR, we conducted experiments on the VisDrone dataset and compared its performance against representative state-of-the-art YOLO series models, FFCA-YOLO, and its baseline RT-DETR-R18. The results are summarized in Table 4.6, where the best values in each column have been highlighted in bold for clarity.
Comparison Experiments for TinyDef-DETR in VisDrone
@@
YOLO10n(Wang et al., 2025) $2.3$ $6.5$ $0.261$ $0.142$ $0.063$ $0.224$
YOLO10s $7.2$ $21.4$ $0.323$ $0.179$ $0.086$ $0.278$
YOLO10m $15.3$ $58.9$ $0.345$ $0.195$ $0.097$ $0.300$
YOLO11n(Khanam and Hussain, 2024) $2.6$ $6.3$ $0.258$ $0.142$ $0.058$ $0.225$
YOLO11s $9.4$ $21.3$ $0.313$ $0.176$ $0.080$ $0.272$
YOLO11m $20.0$ $67.7$ $0.350$ $0.203$ $0.098$ $0.312$
YOLO12n(Tian et al., 2025) $2.6$ $6.3$ $0.259$ $0.142$ $0.057$ $0.224$
YOLO12s $9.2$ $21.2$ $0.312$ $0.176$ $0.081$ $0.274$
YOLO12m $20.1$ $67.2$ $0.336$ $0.192$ $0.094$ $0.298$
YOLO13n(Lei et al., 2025) $2.5$ $6.2$ $0.244$ $0.133$ $0.055$ $0.210$
YOLO13s $9.0$ $20.1$ $0.297$ $0.167$ $0.077$ $0.258$
@@ 5 Conclusions
Conceptualization, J.C. and S.Z.; methodology, J.C.; software, J.C.; validation, J.C. and F.S.; formal analysis, J.C.; investigation, J.C. and W.L.; resources, F.S. and S.Z.; data curation, J.C. and R.Q.; writing—original draft preparation, J.C.; writing—review and editing, S.Z. and F.S.; visualization, J.C.; supervision, S.Z.; project administration, S.Z.; funding acquisition, S.Z. All authors have read and agreed to the published version of the manuscript.

### v5

# Transmission lines;Defect detection;UAV imagery;Small Object Detection;Transformer-based detectors

@@ Abstract
Transmission lines;Defect detection;UAV imagery;Small Object Detection;Transformer-based detectors
@@ 1 INTRODUCTION
With the rapid advancement of smart grid development, transmission line defect detection has increasingly adopted deep learning-based object detection methodsLi et al. (2023). Alongside this progress, numerous general-purpose detectors—such as YOLO(Redmon et al., 2016), DINOZhang et al. (2022), and RT-DETR(Zhao et al., 2024)—have been developed. While these models achieve remarkable success in generic scenarios, most of them on the COCO datasetLin et al. (2014), they face notable limitations in transmission line defect detection under complex conditions involving occlusion, background clutter, and variable lightingLiu et al. (2023). Common issues include missed detections, false positives, and false negatives, which substantially undermine the practical utility of UAV-acquired imagery in real-world transmission line maintenanceDeng et al. (2024). These shortcomings are consistently reported in previous studies and further confirmed by our preliminary experiments, highlighting the critical need for specialized detection frameworks capable of accurately identifying small-scale defects in UAV operational scenarios.
@@
    We propose a lightweight attention block with a dual-domain FSCA (Frequency-Spatial-Channel Attention) mechanism, integrated in a cross-stage partial manner. This design jointly enhances global semantic reasoning and local detail preservation, enabling the network to better capture anisotropic defect patterns such as missing bolts, loose clamps, and sagging wires.
@@ 3.1 Overview
In this study, we employ RT-DETR-r18Zhao et al. (2024) as the baseline detection framework owing to its superior accuracy and robustness in complex scenarios. Traditional DETR models often suffer from slow convergence and substantial computational overhead, which hinder their deployment in UAV-based inspection tasks.
Addressing these inherent limitations, RT-DETR incorporates an enhanced encoder–decoder architecture together with refined query interaction mechanisms, while for the backbone network, RT-DETR-r18 adopts a CNN-based design, namely the popular ResNet-18He et al. (2016). From the perspective of real-time performance, employing a CNN architecture for feature extraction further enhances the real-time capability of DETR-like models, thereby improving their practical applicability. As a result, RT-DETR not only accelerates convergence and improves detection precision, but also preserves the global context modeling capability of Transformer-based frameworks(Vaswani et al., 2017).In addition, its lightweight optimization strategies—specifically, the design choice to apply the encoder only to high-level semantic features (S5) with relatively small spatial dimensions—effectively reduce resource consumption. This balances performance and speed by significantly decreasing computational overhead and accelerating inference without compromising accuracy, thereby making RT-DETR particularly suitable for post-flight UAV imagery analysis, where large volumes of data must be processed efficiently under computational and energy constraints(Tian et al., 2025).
@@
First, the backbone adopts the Edge-Enhanced ResNet (EE-ResNet) structure to strengthen low-level detail preservation. Second, a compact Space-to-Depth Convolution (SPD) module is inserted before each downsampling layer to reorganize spatial features in a manner that preserves critical spatial information while reducing resolution, thereby improving the recall rate of extremely small objects. Third, a lightweight Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM) is embedded into the backbone network to enhance directional sensitivity and fine-grained structural modeling, effectively capturing anisotropic features of typical defects such as missing bolts, loose clamps, or sagging wires. Finally, a Focaler-Wise-SIoU loss function is employed to jointly optimize classification and localization. This loss function adaptively balances easy and hard samples while providing more accurate bounding box regression through shape-aware optimization, which further improves detection robustness for small and ambiguous objects. Detailed descriptions of the proposed modules can be found in Sections 3.2 to  3.5.
@@ 3.2 Edge Enhanced ResNet
Thus, during deployment, EEConv degenerates into a single $3\times 3$ convolution followed by activation:
@@ 4.2 Model Training and Evaluation Metrics
We implement the proposed model on the PyTorch framework and build it upon the Ultralytics RT-DETR backbone initialized with pretrained weights. We conduct training on a workstation equipped with an NVIDIA GeForce RTX 3090 GPU (24 GB) running Ubuntu 20.04, Python 3.10.14, PyTorch 2.3.1, and CUDA 12.1. We train the model for 100 epochs with a batch size of 4 and an input resolution of $640\times 640$, using the AdamW optimizer with an initial learning rate of $1.0\times 10^{-4}$, a final learning rate factor $lrf=1.0$, momentum of 0.9, and weight decay of $1.0\times 10^{-4}$. We apply a warm-up of 2000 iterations with momentum set to 0.8 and bias learning rate set to 0.1. For evaluation, we use the validation split following the COCO protocol with an IoU threshold of 0.7 and a maximum of 300 detections per image.
@@ 4.4 Ablation Study of the Proposed Method
[…]
@@
Cross-Stage Dual-Domain Multi-Scale Attention Module (CSDMAM) The introduction of CSDMAM demonstrates its effectiveness in enhancing the model’s capability to capture small-scale defects under complex backgrounds. Although an intermediate configuration without the loss function shows minor fluctuations (e.g., precision 0.425 vs. 0.417, recall 0.190 vs. 0.208), the overall performance trend confirms the complementary role of CSDMAM when integrated with other modules. Notably, when CSDMAM is combined with both EER and SPD, the $\mathrm{AP}{s}$ improves markedly from 0.083 to 0.119, highlighting its strength in boosting sensitivity to small objects. This suggests that CSDMAM effectively suppresses background interference while enhancing multi-scale feature representation, which is crucial for distinguishing subtle defects from visually similar structures in transmission line imagery. The observed gains in $\mathrm{AP}{s}$ validate the necessity of CSDMAM for robust small-target detection, particularly when used synergistically with the proposed loss function.
@@ 4.5 Comparisons With Previous Methods
Table LABEL:tab:model_performance summarizes the overall performance of representative detectors on the CSG-ADCD dataset. The comparison covers lightweight and standard variants of YOLO (v5, v8, 10, 11), Transformer-based models such as RT-DETR and Conditional-DETR, as well as classical baselines including RetinaNet and DETR. Key indicators, including the number of parameters, computational complexity (GFLOPs), Precision, Recall, $\text{mAP}_{50}$, and $\text{mAP}_{50:95}$, are reported to provide a comprehensive view of accuracy–efficiency trade-offs. The best values in each column are highlighted in bold for clarity.
For a fair comparison, all methods are evaluated under a unified protocol—an input resolution of $640\times 640$, identical NMS settings, and the same maximum number of detections per image.
@@
Table 3: Model Performance Evaluation: Precision, Recall, and mAP for CSG-ADCD
@@
YOLO v5n $2.2$ $5.8$ 0.707 $0.131$ $0.131$ 0.062 80.062\,8
YOLO v5s $9.1$ $23.8$ $0.346$ $0.148$ $0.140$ 0.064 60.064\,6
YOLO v5m $25.1$ $64.0$ $0.548$ $0.164$ $0.162$ 0.071 70.071\,7
YOLO v8n $2.7$ $6.8$ $0.580$ $0.134$ $0.133$ 0.061 30.061\,3
YOLO v8s $11.1$ $28.5$ $0.375$ $0.155$ $0.147$ 0.066 10.066\,1
YOLO v8m $23.2$ $67.5$ $0.403$ $0.169$ $0.163$ 0.071 90.071\,9
YOLO 10n(Wang et al., 2025) $2.7$ $8.2$ 0.707 $0.125$ $0.133$ 0.062 70.062\,7
YOLO 10s $8.0$ $24.5$ $0.536$ $0.144$ $0.157$ 0.072 50.072\,5
YOLO 10m $16.6$ $63.5$ $0.386$ $0.170$ $0.162$ 0.075 30.075\,3
YOLO 11n(Khanam and Hussain, 2024) $2.6$ $6.3$ $0.638$ $0.135$ $0.136$ 0.064 10.064\,1
YOLO 11s $9.4$ $21.3$ $0.516$ $0.166$ $0.155$ 0.071 70.071\,7
YOLO 11m $20.0$ $67.7$ $0.399$ $0.181$ $0.173$ 0.077 60.077\,6
RT-DETR-l $32.8$ $108.0$ $0.480$ $0.189$ $0.199$ 0.089 00.089\,0
RT-DETR-x $65.5$ $222.5$ $0.485$ $0.199$ $0.216$ 0.099 00.099\,0
RT-DETR-R18 $19.8$ $57.0$ $0.369$ $0.177$ $0.163$ 0.073 80.073\,8
RT-DETR-R34 $31.2$ $88.8$ $0.354$ $0.127$ $0.117$ 0.053 30.053\,3
RT-DETR-R50 $42.9$ $134.8$ $0.363$ $0.120$ $0.115$ 0.051 40.051\,4
RT-DETR-R101 $74.7$ $247.1$ $0.467$ $0.185$ $0.193$ 0.087 90.087\,9
Conditional-DETR-R18(Meng et al., 2023) $0.309$ $0.068$ $0.038$ 0.011 00.011\,0
Conditional-DETR-R34 $0.447$ $0.082$ $0.034$ 0.009 00.009\,0
Conditional-DETR-R50 $44.0$ $89.5$ $0.568$ $0.159$ $0.172$ 0.066 00.066\,0
DETR-R50(Carion et al., 2020) $0.301$ $0.138$ $0.124$ 0.043 00.043\,0
RetinaNet(Lin et al., 2018) $12.0$ $33.8$ $0.394$ $0.029$ $0.013$ 0.004 00.004\,0
DINO(Zhang et al., 2022) $0.496$ $0.257$ $0.224$ 0.098 60.098\,6
FFCA-YOLO(Zhang et al., 2024) $7.1$ $51.4$ $0.680$ $0.130$ $0.142$ 0.059 60.059\,6
@@
To further highlight performance on small objects and defect-specific categories, Table LABEL:tab:model_performance2 reports additional evaluation metrics. These include Recall and $\text{mAP}_{50}$ for Polluted Insulator (jyzwh) and Damaged Pole (dgss), alongside scale-sensitive measures $\text{AP}_{s}$ and $\text{AP}_{m}$. Such results enable a fine-grained comparison of detection
@@ 4.6 Generalization Experiments for the Proposed Methods
Since small objects account for a large proportion in our dataset, we further assess the generalization ability of the proposed method on the VisDrone dataset (Zhu et al., 2022). VisDrone is a large-scale benchmark specifically constructed for UAV-based visual perception, created by the Aerial Vision Group at the University of Science and Technology of China. It contains over 26,000 high-resolution images and 263 video sequences captured across diverse urban and suburban scenes (streets, rooftops, playgrounds, traffic), with more than 100,000 manually annotated instances covering 10 common categories (e.g., pedestrian, car, bus, truck, bicycle). Each instance is labeled with bounding box, class, occlusion level, and truncation ratio. Owing to dense object distributions, the high proportion of small objects, severe occlusion, and large variations in viewpoint and illumination, VisDrone poses considerable challenges for detection and tracking models and is widely used to evaluate robustness and generalization in real-world UAV scenarios.
For a fair comparison, all methods are evaluated under the same protocol—an input resolution of $640\times 640$, identical NMS settings, and the same maximum number of detections per image. We compare TinyDef-DETR against representative state-of-the-art YOLO-series models, FFCA-YOLO, and its baseline RT-DETR-R18. The results are summarized in Table 5, where the best value in each column is highlighted in bold.
Table 5: Comparison Experiments for TinyDef-DETR on VisDrone
@@
YOLO v10n(Wang et al., 2025) $2.3$ $6.5$ $0.261$ $0.142$ $0.063$ $0.224$
YOLO v10s $7.2$ $21.4$ $0.323$ $0.179$ $0.086$ $0.278$
YOLO v10m $15.3$ $58.9$ $0.345$ $0.195$ $0.097$ $0.300$
YOLO v11n(Khanam and Hussain, 2024) $2.6$ $6.3$ $0.258$ $0.142$ $0.058$ $0.225$
YOLO v11s $9.4$ $21.3$ $0.313$ $0.176$ $0.080$ $0.272$
YOLO v11m $20.0$ $67.7$ $0.350$ $0.203$ $0.098$ $0.312$
YOLO v12n(Tian et al., 2025) $2.6$ $6.3$ $0.259$ $0.142$ $0.057$ $0.224$
YOLO v12s $9.2$ $21.2$ $0.312$ $0.176$ $0.081$ $0.274$
YOLO v12m $20.1$ $67.2$ $0.336$ $0.192$ $0.094$ $0.298$
YOLO v13n(Lei et al., 2025) $2.5$ $6.2$ $0.244$ $0.133$ $0.055$ $0.210$
YOLO v13s $9.0$ $20.1$ $0.297$ $0.167$ $0.077$ $0.258$
@@ 5 Conclusions
Conceptualization, J.C. and S.Z.; methodology, J.C.; software, J.C.; validation, J.C. and F.S.; formal analysis, J.C.; investigation, J.C. and W.L.; resources, F.S. and S.Z.; data curation, J.C.; writing—original draft preparation, J.C.; writing—review and editing, S.Z. and F.S.; visualization, J.C.; supervision, S.Z.; project administration, S.Z.; funding acquisition, S.Z. All authors have read and agreed to the published version of the manuscript.